How many gallons of 30% alcohol solution and 75% alcohol solution must be mixed to get 3 gallons of 45% alcohol solution?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol type I 30 ---------------- x gallons
alcohol type II 75 ---------------- 3 - x gallons
Mixture 45.00% ---------------- 3

30x+75(3-x)=45* 3

30x+225 -75x=135
30x-75x =135-225
-45x=-90
/ -45
x =2 gallons 30.00% Alcohol type I
1 gallons 75.00% alcohol type II